# Making congestion control robust to per-packet load balancing in datacenters

### v1

@@ Abstract
Per-packet load-balancing approaches are increasingly deployed in datacenter networks. However, their combination with existing congestion control algorithms (CCAs) may lead to poor performance, and even state-of-the-art CCAs can collapse due to duplicate ACKs. A typical approach to handle this collapse is to make CCAs resilient to duplicate ACKs.
In this paper, we first model the throughput collapse of a wide array of CCAs when some of the paths are congested. We show that addressing duplicate ACKs is insufficient. Instead, we explain that since CCAs are typically designed for single-path routing, their estimation function focuses on the latest feedback and mishandles feedback that reflects multiple paths. We propose to use a median feedback that is more robust to the varying signals that come with multiple paths. We introduce MSwift, which applies this principle to make Google’s Swift robust to multi-path routing while keeping its incast tolerance and single-path performance. Finally, we demonstrate that MSwift improves the 99th-percentile FCT by up to 25%, both with random packet spraying and adaptive routing.
@@ I Introduction
Motivation. In recent years, datacenter networks have faced a dual challenge: a dramatic increase in traffic volumes alongside client demands for shorter job completion times. This combination has sparked widespread calls for better utilization of existing infrastructure. A prominent approach involves per-packet load balancing, including adaptive routing (AR) [1, 2, 3], e.g., in NVIDIA’s Spectrum-X [4, 5]; random packet spraying (RPS) [6, 7, 8]; and adaptive packet spraying algorithms [9, 10], e.g., in the Ultra Ethernet Consortium specification [11].
However, these per-packet load-balancing approaches are particularly challenging for state-of-the-art datacenter congestion control algorithms (CCAs) that operate with single-path assumptions, such as DCTCP, which is deployed in Meta’s production datacenters [12, 13, 14], and Swift, which is deployed in Google’s [15]. The throughput of these CCAs simply collapses when using such per-packet load balancing. Long-term congestion on one of the paths due to single-path background elephant flows or asymmetric topologies only worsens this collapse.
To handle this mismatch, (1) one alternative solution in the literature is to limit the per-packet load balancing policy. For example, Hermes [16] reduces path changes to avoid getting congestion signals from multiple routes. (2) Another approach is to focus only on the packet reordering problem, especially with TCP [6, 7, 17, 18, 19, 20, 21, 22, 23, 24]. When a packet traverses a high-latency path and arrives later than its predecessors, it can trigger duplicate ACKs and cause unnecessarily aggressive multiplicative decreases in the congestion window. Similarly, since many datacenter CCAs like DCTCP and Swift also rely on packet reordering as a congestion signal, having a reordering buffer to solve the reordering problem can help them as well.
@@
Unfortunately, per-packet load balancing affects nearly all congestion signals, not only reordering. Thus, to avoid throughput collapse in modern CCAs that rely on several congestion signals, dealing only with reordering is insufficient, and the other congestion signals should be handled as well. Otherwise, if a single path is congested out of $n$ paths, then packets traversing the congested path will cause the sender to reduce its rate, even if all other paths are uncongested. This overreaction can lead to substantial under-utilization of available network capacity and to larger flow completion times (FCTs). The above congested path can affect about any congestion signal, e.g., (a) a large delay in delay-based CCAs such as Swift; (b) an ECN-marked packet in ECN-based CCAs such as DCTCP; or (c) a congestion signal in In-band Network Telemetry (INT)-based CCAs such as Poseidon [25].
In this paper, we argue that there is a need for an approach that is general enough to cover these various congestion signals within the CCA given any per-packet load balancing. We propose a median-based framework that is robust enough to address these congestion signals. We further delve into Swift to produce a specific implementation example that can be practically evaluated.
@@
Contributions. We introduce a fundamental model of throughput collapse induced by per-packet load-balancing. We assume a simple round-robin packet spraying scenario, where packets go through a congested path with probability $q$, while other paths stay uncongested, and exhibit a general rule for additive-increase multiplicative-decrease (AIMD) flows: Throughput collapses as $\Theta(1/\sqrt{q})$. We show that it applies to the (1) TCP, (2) DCTCP, and (3) Swift reordering-sensitive CCAs. We also demonstrate that the same asymptotic rule applies to two reordering-insensitive CCAs: (4) FastFlow [26], which uses additive increase and both additive and multiplicative decrease, and relies on delay and ECN congestion signals, and (5) a reordering-resilient Swift variant, which effectively only uses delay. This proves that only solving reordering is insufficient.
We further validate the Swift and reordering-resilient Swift models through simulations and observe a good fit between models and simulations, using both random and round-robin packet-spraying simulations.
To solve the per-packet multi-path collapse, we propose a median-based framework that tracks recent congestion signals and responds based on their median value. This simple approach applies to a wide set of signals, and remains robust against varying per-packet multi-path signals.
We focus our implementation and evaluation on Swift. We suggest and implement MSwift, where we apply our median-based framework to a reordering-resilient Swift baseline. We also analyze several approaches for the history sizes to be used in the median computation, ultimately proposing a bounded-window one. We examine several multi-path scenarios: permutation scenarios where MSwift is either alone in the network or also suffers from the interference of single-path elephant background flows, and incast scenarios. Extensive evaluations using 3-layer fat-tree topologies show that MSwift improves the mean throughput and 99th-percentile FCT by up to 25% in permutation scenarios compared to our baseline, using both the RPS and AR per-packet routing schemes. We also show that MSwift retains our baseline’s tolerance to incast traffic, and retains Swift’s performance in single-path scenarios.
In summary, our main contributions are: (1) the fundamental $\Theta(1/\sqrt{q})$ throughput-collapse rule for a large class of CCAs; and (2) the median-based framework solution to fight per-packet multi-path throughput collapse.
We intend to release our code on Github upon publication.
Related Work. A major obstacle to per-packet load balancing is that different paths may experience varying latencies, resulting in packet reordering that triggers congestion avoidance mechanisms and reduces throughput. This phenomenon has been extensively studied for TCP, with numerous solutions proposed over the years. Leung et al. [18] provide an overview of those TCP reordering solutions and discuss them in detail. A first approach [24] either temporarily disables congestion control or performs recovery during congestion avoidance upon reordering. A common technique [20, 21] is to readjust the duplicate ACK threshold. Another mechanism [22] is to delay the congestion response for one RTT after receiving the first duplicate ACK. It is also possible to readjust RTO timer calculations [23]. More recently, it has been suggested to use network coding [19]. Another approach is to completely ignore reordering as a loss signal and rely solely on an RTO timer with adjusted parameters [17]. Finally, Lazy TCP [7] was specifically designed for datacenters. It waits for reordered packets and decreases the window only upon consecutive packet reordering. Unfortunately, these solutions do not provide solutions for congestion signals beyond reordering.
Congestion signals in the literature include ECN, delay, and INT signals. For example, DCTCP [12] uses both loss signals and ECN marking. Swift [15] employs loss signals and incorporates the delay of the last received packet. FastFlow [26] uses only packet trimming and timeouts as loss signals, ignoring reordering altogether, additionally relying on delay and ECN feedback. Poseidon [25] mainly relies on INT signals representing the max per-hop delay (MPD) and on loss signals. DCQCN [27] relies on ECN marking, relayed back through congestion notification packets (CNPs), together with hop-by-hop, link-layer Priority-based Flow Control (PFC).
Load balancing solutions such as Hermes [16] try to partially avoid this mismatch of CCA and per-packet load-balancing interaction by limiting the per-packet load balancing rerouting policy. Instead, we want to keep it unchanged.
@@ II The $\Theta(1/\sqrt{q})$ CCA Multi-Path Rule
II-A Outline and assumptions.
Figure 1: Periodic round-robin scenario with one congested path.
Overview. In this section, we model the multi-path performance of a single flow. As Fig. 1 illustrates, it employs round-robin packet spraying across $n$ paths. We assume that a single path out of $n$ experiences congestion, i.e., with probability $q=1/n$. We also assume a constant-delay congestion to make the model tractable. We denote the large congested RTT as $T_{l}$ and the short uncongested RTT as $T_{s}$.
Our analysis reveals that if the flow uses TCP, DCTCP, Swift, FastFlow, or a reordering-resilient Swift as CCAs, it will experience a throughput collapse that follows a common scaling rule:
@@
Table I and Fig. 2 summarize our theoretical models. [^1] To plot Fig. 2, we assume that Swift uses parameters $ai=1$, $mm\equiv max\_mdf=0.5$, $\beta=0.8$ (default values in the htsim simulator [28]), $base\_target+topology\_scaling=21\mu s$ and htsim default flow-based scaling parameters, that FastFlow uses the paper’s default parameters $targetRTT=1.5\cdot baseRTT$, $mi=2$, $md=2$, $fi=0.25$, $fd=0.8$ and use $\alpha=\frac{T_{s}}{baseRTT}=1.2$. We assume that congestion increases the RTT to be $T_{l}=2.5T_{s}$. We assumed $T_{s}=14.4\mu s$ so that $baseRTT=12\mu s$ for FastFlow to fit the default values. The takeaway is that the impact of even a few congested paths is rather dramatic. Having 2% of congested paths rather than 1% decreases the throughput by a factor of $\sqrt{2}\approx 1.4\times$, an awful impact in datacenters, with an even worse effect on 99th-percentile FCTs.
TABLE I: CCA multi-path throughput models when congestion occurs with probability $q$ and leads to a path delay of $T_{l}$ rather than $T_{s}$. MSS stands for maximum segment size, $mm\equiv max\_mdf$ and aiai are Swift parameters. The throughput of all these CCAs falls like $\Theta(1/\sqrt{q})$, experiencing a sharp drop even with little congestion.
Thm. CCA Throughput Model
Eq. 1 TCP $\frac{MSS}{T_{s}}\frac{1.22}{\sqrt{q}}$
Eq. 3 DCTCP (upper bound) $\frac{MSS}{T_{s}}\frac{1.22}{\sqrt{q}}$
Eq. 4 Swift $\frac{MSS}{T_{s}}\frac{\sqrt{\left(1/mm-1/2\right)\cdot ai}}{\sqrt{q}}$
Eq. 5 FastFlow $\Theta(1/\sqrt{q})$
Eq. 6 Reordering-resilient Swift $\Theta(1/\sqrt{q})$
Figure 2: Model comparison using the theoretical results in Table I. Under all considered CCAs, throughput in round-robin packet spraying collapses when one path out of $n$ is congested, i.e., congestion probability is $q=1/n$.
Intuition. As Fig. 3 illustrates for Swift, our analysis reveals the intuition behind this rule for algorithms employing an additive increase and multiplicative decrease (AIMD) policy for their congestion window $W$. Intuitively, they send a number of packets proportional to $W^{2}$ before congestion happens, which occurs every $n$ packets, so $W^{2}\propto n=1/q$ and $W\propto 1/\sqrt{q}$. The rule is closely related to the Mathis $\Theta(1/\sqrt{p})$ throughput model for TCP given a loss rate of $p$ [29]. The reason is that in TCP, both losses and late packets are detected through the same mechanism of three duplicate ACKs. But both rules are not identical. If path delays in the congested path were replaced by drops, as in the Mathis model, the $\Theta(1/\sqrt{p})$ model would not hold for several of our CCAs given some topology parameters. In fact, FastFlow and reordering-resilient Swift sometimes do not even reduce their congestion windows upon packet drop.
Mathis et al. [29] assumptions. We follow the Mathis model assumptions that the flow avoids retransmission timeouts, always has sufficient receiver window and data to send, the receiver is acknowledging every packet without delay, multiple congestion signals within one round trip are treated as a single one, and we can neglect rounding of the congestion window (cwnd) upon reduction, cwnd limits, and details of data recovery and retransmission.
We supplement these with additional assumptions to be able to derive a closed-form model. We assume that the buffer size is large enough that the loss probability is negligible, that the congestion window is sufficiently large that (1) after every packet traversing the long path, three packets from short paths return first, i.e., congestion causes three duplicate ACKs, and (2) the congestion window does not stall while waiting for ACKs of long-path packets. We also assume that $n/W$ is large enough that multiple window increases occur between consecutive decreases. We further assume for all models that all packet sizes are the maximum segment size (MSS).
II-B Reordering-sensitive models.
TCP model. We first model the behavior of several reordering-sensitive CCAs, starting with TCP.
@@
Under the stated assumptions, the throughput of a single TCP flow is
@@
Proof:
DCTCP model. For this model of DCTCP [12], we assume that congestion is localized at one buffer, and distinguish the result based on whether the added long-path congestion delay ($T_{l}-T_{s}$) is below the ECN buffer threshold (our default assumption) or not. Note that the ECN threshold is often set around 1 BDP [30]. Both cases yield an $O\left(1/\sqrt{q}\right)$ throughput, with the same main conclusions on throughput collapse.

### v2

# Congestion Control for Spraying with Congested Paths

@@ Abstract
Packet spraying approaches are increasingly deployed in datacenter networks. However, their combination with existing congestion control algorithms (CCAs) may lead to poor QoS, especially when some of the paths are congested.
In this paper, we first model the throughput collapse of a wide array of CCAs when some of the paths are congested. We explain that since CCAs are typically designed for single-path routing, their estimation function focuses on the latest feedback and mishandles feedback that reflects multiple paths. We propose using a median feedback that is more robust to the varying signals that come with multiple paths. We introduce MSwift and MNSCC, which apply this median principle to Google’s Swift and Ultra Ethernet’s NSCC. We demonstrate that they can improve both CCAs, reaching better QoS both under congested paths and in uncongested networks.
@@ I Introduction
RoCEv2 (RDMA over Converged Ethernet version 2) is used in many datacenter networks. It enables a wide set of applications, from AI training to Storage Area Networks (SANs). Unfortunately, many existing RoCE NIC implementations have historically been sensitive to reordering due to simple Go-Back-N (GBN) loss-recovery designs [1]. Recent NICs support out-of-order delivery, but only for a subset of RDMA operations such as RDMA WRITE [2]. However, other operations might still suffer from poor performance due to packet reordering. For example, RDMA READ operations send a multi-packet response where packets need to arrive in order, or else the late packets will be discarded and will have to be resent [1]. As a result, RoCEv2 datacenters often rely on ECMP-based load-balancing techniques with a single path per flow. These techniques either prevent reordering entirely, as in Meta’s enhanced ECMP [3], or minimize it by using ECMP but allowing occasional re-pathing, as in FlowBender and Google’s PLB [4, 5].
However, with modern AI training workloads, such techniques are prone to flow collisions and yield poor performance [6, 7, 8]. Thus, many companies increasingly rely on per-packet load balancing, also called packet spraying, to efficiently use their networks. For example, Alibaba implements host-based Oblivious Packet Spraying (OPS) [9, 10, 11, 6]; NVIDIA’s Spectrum-X [2, 12] implements switch-based Adaptive Routing (AR) [13, 7, 14]; and the recent Ultra Ethernet Consortium (UEC) specification [15] recommends host-based adaptive packet spraying algorithms such as REPS and STrack [16, 17].
Therefore, modern networks must be efficient for the cases where per-packet spraying and per-flow ECMP variants must coexist. For example, NVMe-over-Fabric servers [18] heavily rely on RDMA READ operations which cannot run efficiently with per-packet spraying, whereas the NVIDIA NCCL collective library predominantly employs RDMA WRITE specifically to allow spraying [19, 8]. Thus, most switch vendors support a concurrent use of both types of load balancing on the same network by marking for eligibility packets that can be sprayed (e.g., Cisco [20] and NVIDIA [2]). In this case, the network congestion is highly non-uniform: some links are much more congested, as ECMP flows and sprayed packets are using these same links.
The Congestion Control Algorithm (CCA) of the sprayed flows suffers in two ways from these congested links. The first problem is well known: the asymmetry causes packet reordering. This has been well studied in the past, especially with TCP [9, 10, 21, 22, 23, 24, 25, 26, 27, 28]. When a packet traverses a high-latency path and arrives later than its predecessors, it can trigger duplicate ACKs and cause unnecessarily aggressive multiplicative decreases in the congestion window. This reordering problem can be mitigated by using Selective ACK (SACK) and delaying the time to perform congestion avoidance [10].
@@
In this paper, we identify a second, previously unknown problem: a throughput collapse due to the recurrent congestion signals from the congested links. Even if very few links are congested, packets crossing the congested links will repeatedly come back with congestion signals and cause the sender to aggressively reduce its rate. This overreaction can lead to substantial under-utilization of available network capacity and to increased Collective Completion Times (CCTs) in AI training. The congested path can affect about any congestion signal, e.g., (a) a large delay in delay-based CCAs such as Google’s Swift [29]; (b) an ECN-marked packet in ECN-based CCAs such as DCTCP, which has been deployed in Meta’s production datacenters [30, 31, 32]; or (c) a congestion signal in In-band Network Telemetry (INT)-based CCAs such as Poseidon [33].
We could think of several natural solutions to this throughput collapse. (i) A first approach would be to allocate a different congestion window for each path, as in MPTCP or MP-RDMA [34]. However, subflow-based load-balancing exhibits poor performance [35]. Packet spraying techniques employ a large number of paths for efficient load-balancing. The resulting large number of windows would be hard to implement at scale, and many tiny windows would achieve poor results. Furthermore, in switch-based spraying techniques like AR, the hosts do not even know the path of the packets. (ii) A second approach would be to employ packet spraying techniques like REPS that avoid sending over congested links. We show in evaluations that this does not help sufficiently (Section IV). (iii) A third approach would be to separate ECMP flows from sprayed packets using traffic classes, and provide a fair scheduling in switches. In addition to its configuration complexity and interference with the existing QoS policies, we show that this approach still suffers from the throughput collapse, since sprayed packets keep encountering more congestion at shared links, and therefore hosts keep receiving recurrent congestion signals (Section IV).
@@
Contributions. In this paper, we make three main contributions. First, we introduce a fundamental model of throughput collapse induced by the joint use of packet spraying and ECMP. We assume a simple scenario where packets go through a congested path with probability $q$, while other paths stay uncongested. We demonstrate a general rule: Throughput collapses as $\Theta(1/\sqrt{q})$. We prove that it applies to five different CCAs: (1) TCP (New Reno); (2) DCTCP; (3) Swift; (4) SMaRTT/FastFlow [36], the basis of the NSCC CCA in the UEC specification [15] that is designed for packet spraying; and (5) a reordering-resilient Swift variant. This occurs even though these CCAs use different congestion signals in different ways.
Second, to solve the throughput collapse, we propose to adapt CCAs by using a general median-based framework. Instead of responding only to the latest congestion signal, as we would when signals come from a single path, this simple approach tracks recent congestion signals and reacts based on their median value. Thus, this approach remains robust against varying per-packet multi-path signals, while following the intended goals of the CCA. We explain how this framework can also be fine-tuned using a Nyquist perspective. We illustrate the framework with two CCAs: (i) based on Google’s Swift, we implement MSwift, where we apply our median-based framework to a reordering-resilient Swift baseline; and (ii) based on UEC’s NSCC, we introduce MNSCC. Extensive evaluations using 3-layer fat-tree topologies show that MSwift improves the sprayed-traffic CCT inflation in the presence of ECMP flows by up to $8.6\times$ using REPS. It outperforms both Swift and NSCC in all 3 compared spraying algorithms: OPS, AR and REPS. In addition, using fair DRR scheduling at switches exhibits similar results.
Our third contribution is to show that this new CCA framework does not only help with permanent ECMP congestion, but also with temporary congestion without ECMP flows. We simulate training a Llama 70B model using HSDP, and find that MSwift outperforms all other algorithms on all spraying techniques. Its CCT inflation with REPS is $1.8\times$ smaller than a reordering-resilient Swift and $2.3\times$ smaller than NSCC.
@@ II The $\Theta(1/\sqrt{q})$ CCA Multi-Path Rule
Figure 1: Scenario with single congested path.
In this section, we start by presenting a theoretical analysis that gives us a more rigorous understanding of the impact of a few congested paths on packet-sprayed flows using existing CCAs. We show that it causes the throughput collapse of these flows. This result leads us to propose MSwift in Section III.
II-A CCAs
We model and demonstrate throughput collapse in five core CCAs that can be used in datacenters and rely on different congestion signals.
(1) TCP. We start with the basic TCP New Reno, as it is a fundamental inspiration for later CCAs.
(2) DCTCP, which uses both loss signals and ECN marking.
(3) Swift. Google’s Swift [29] strongly relies on packet delays, as well as on loss signals.
The above 3 CCAs are reordering-sensitive. We also demonstrate that the same throughput collapse applies to two additional reordering-insensitive CCAs:
(4) SMaRTT. SMaRTT [36], also known as FastFlow, is the initial default CCA for REPS. It is important as it forms the basis for NSCC, the mandatory CCA in the UEC specification [15]. It relies on delay and ECN congestion signals.
(5) Reordering-resilient Swift. This variant is adopted to neutralize the impact of reordering and show that throughput collapse happens despite it.
II-B Model
Fig. 1 illustrates our theoretical model. We consider a single flow that sprays packets across $n$ paths. We assume that a single path out of $n$ always experiences congestion. Therefore, the probability of a random packet to experience congestion is $q=1/n$. We also assume a constant-delay congestion to make the model tractable. We denote the larger RTT of the congested path as $T_{l}$ and the shorter RTT of the uncongested paths as $T_{s}$.
II-C Results
Our analysis reveals that if the flow uses TCP, DCTCP, Swift, SMaRTT, or a reordering-resilient Swift as CCAs, it will experience a throughput collapse that follows a common scaling rule:
@@ II-C Results
TABLE I: CCA multi-path throughput models when congestion occurs with probability $q$ and leads to a path delay of $T_{l}$ rather than $T_{s}$. MSS stands for maximum segment size, $mm\equiv max\_mdf$ and $ai$ are Swift parameters. The throughput of all these CCAs falls like $\Theta(1/\sqrt{q})$, experiencing a sharp drop even with little congestion.
Thm. CCA Throughput Model Collapse Reason
Thm. 1 TCP $\frac{MSS}{T_{s}}\frac{1.22}{\sqrt{q}}$ Packet Reordering
Eq. 3 DCTCP (upper bound) $\frac{MSS}{T_{s}}\frac{1.22}{\sqrt{q}}$ Packet Reordering
Eq. 4 Swift $\frac{MSS}{T_{s}}\frac{\sqrt{\left(1/mm-1/2\right)\cdot ai}}{\sqrt{q}}$ Packet Reordering
Eq. 5 SMaRTT $\Theta(1/\sqrt{q})$ Delay + ECN Signals
Eq. 6 Reordering-resilient Swift $\Theta(1/\sqrt{q})$ Delay Signal
Given the general model above, we make additional simplifying assumptions that mostly follow the standard Mathis model assumptions [37]. For example, we assume that paths are taken in a round-robin order rather than at random, and that the flow avoids retransmission timeouts. We also make CCA-specific assumptions. For example, for DCTCP, we distinguish whether the long path triggers congestion or not. All assumptions and proofs are detailed in Appendix A. Under these assumptions, as Table I summarizes, we obtain the following theorems:
@@
The throughput of a single TCP flow is